Use this information for the following questions.

  • 40 year old male
  • resting heart rate 65 beats/minute
  • maximal heart rate 195 beats/minute
  • weight 180 pounds
  • VO2max55 ml/kg/min
  • Treadmill speed = 3.8 mph
  • Exercise prescription, 65% of VO2max
  • Vo2 ml/kg/min walking equation= 3.5 + (.1 x S) + [1.8*S* (grade/100)]
  1. What is his target heart rate range using an Age Predicted Max Heart Rate Method, using 50-70% of his Age-predicted Max Heart Rate? (2 points)
  1. What is his target heart rate range, using the Modified Karvoven Method/Modified Heart Rate Reserve method, if we want him to exercise at a range 50-70% of his Heart Rate Reserve? (2 points)

  1. What is his target VO2 if you want him to exercise at 65% of VO2max?  (Straight percentage, NOT reserve).  (2 points)
Answer #1

1. The maximum Target heart range by using Age predicted Max Heart Rate method, with the % of 50-70.
Solution;
age predicted max heart rate = 220- age
= 220-40
= 180 bpm

then, the Max target heart range is,
for lower range of 50% = 180 x 50/100 = 90 bpm
for higher range of 70% = 180 x 70/100 = 126 bpm
So the Max Target Heart range would be 90-126 bpm.

2. The Target heart rate range using Modified karvonen method ;

Solution;
given resting HR =65 bpm
age predicted Max HR = 180 bpm
max HR - resting HR = 180-65 = 115

the max target heart range is,
for the lower range of 50% = (115 x 0.5) + 65 = 122
for the lower range of 70% = (115 x 0.7) + 65 = 145

So the target heart range is 122 bpm to 145 bpm.

3. For the Target VO2 with the exercise level of 65% of VO2 Max
Solution,

Target VO2 = exercise level ( its intensity ) x VO2 max
= 0.65 x 55
= 35 ml.kg.mins